# Successive randomized compression: A randomized algorithm for the compressed MPO–MPS product

### v1

@@ Appendix C Implementing SRC with a tolerance
In section 3, we presented a version of SRC where the output bond dimension $\overline{\chi}$ is provided as an input to the algorithm. This section will describe an adaptive implementation of SRC that chooses the bond dimension adaptively to meet a tolerance. Section C.1 discusses error estimation, section C.2 uses the error estimate to implement SRC adaptively, and section C.3 optimizes the adaptive implementation using QR factorization updating algorithms.
@@ C.1.1 Leave-one-out error estimation: Matrix case
Surprisingly, the QR factorization C.1 already contains enough information to compute an estimate of the error $\mathrm{Err}_{p}$. First, compute the inverse-adjoint of the matrix $R$:
@@
That is, $\smash{\widehat{\mathrm{Err}}}^{2}$ is a statistically unbiased estimate for the mean-square error $\mathrm{Err}_{p-1}^{2}$ of the rank-$(p-1)$ randomized QB approximation. Consequently, $\smash{\widehat{\mathrm{Err}}}^{2}$ is an overestimate for $\mathrm{Err}_{p}^{2}$, on average.
@@ C.1.3 Norm estimation
This is an instance of the Girard–Hutchinson norm estimator, closely related to the Girard–Hutchinson trace estimator [38, §4]. In particular, we have the following standard guarantee [38, §4.8]:
@@
Under the hypotheses of Theorem 4, the squared norm estimate $\smash{\widehat{\mathrm{Norm}}}^{2}$ is a statistically unbiased estimate for $\left\|\smash{B}^{(j)}\right\|^{2}_{\rm F}$. That is,
@@ C.2 Adaptive determination of bond dimension
Using the leave-one-out error estimate C.2 and norm estimate C.3, we can adaptively choose the dimension $\overline{\chi}$ adaptively for each bond in the output MPS using a tolerance. Let $\tau_{\rm abs}$ and $\tau_{\rm rel}$ be absolute and relative tolerances, and let $\overline{\chi}_{0}$ and $\Delta_{\chi}$ be a user-specified minimum bond dimension and increment. Begin each step $j$ of SRC using the minimum bond dimension $\overline{\chi}_{0}$, obtaining a $j$th site tensor $\eta^{(j)}$. Then, do the following steps:
@@
    Update randomized QB approximation. Compute the additional columns $\Delta_{\chi}$ of $Y^{(j)}$ using the (possibly updated) intermediate tensor $C^{(j-1)}$. Recompute (or update, see section C.3) the QR factorization of $Y^{(j)}$. Go to step 1.
When implemented using QR updating rather recomputation from scratch, as described in the next section, this adaptive version of the MPO–MPS multiplication algorithm has the same time complexity as reported in section 3.5. In our experiments in section 4.4.3, we use parameters $\overline{\chi}_{0}=2$, $\Delta_{\chi}=3$, and use only a relative tolerance (setting $\tau_{\rm abs}=0$). To implement with oversampling, we set the relative tolerance to be $0.1$ times the requested tolerance and run a final truncation with the requested tolerance.
@@ C.3 Final optimization: Updating the QR factorization
In the adaptive MPO–MPS multiplication algorithm (section C.2), we continue to accumulate “columns” in the 2- or 3-tensor $Y^{(j)}$ until a tolerance is satisfied. Recomputing a QR factorization each time we do this is computationally expensive, and we can make the algorithm more efficient by instead updating the QR factorization to include new columns. This material is standard but perhaps not well-known, so we provide a brief discussion.
@@
Another matrix that must be updated to implement the adaptive MPO–MPS multiplication algorithm efficiently is the matrix $G$ used in error estimation C.2. Again using our superscript-less notation, we begin with $G=R^{-\dagger}$ and with to update $G$ to
@@
We recognize the original $G$ matrix in the top left corner of this matrix. To finish the formula for $G^{\prime}$, we just need to compute $(R^{\prime\prime})^{\dagger}$ and form the matrix triple product $-((R^{\prime\prime})^{-\dagger}(R^{\prime})^{\dagger})G$. Since $R^{\prime\prime}$ is a small $\Delta_{\chi}\times\Delta_{\chi}$ matrix, this is efficient to do.
@@ Appendix D Full operation counts
Full operation counts for the methods in terms of all five parameters $n$, $d$, $\chi$, $D$, and $\overline{\chi}$ are presented in Table 2. For simplified operation counts, see Table 1.

### v2

@@ Appendix C Implementing SRC with a tolerance
In section˜3, we presented a version of SRC where the output bond dimension $\overline{\chi}$ is provided as an input to the algorithm. This section will describe an adaptive implementation of SRC that chooses the bond dimension adaptively to meet a tolerance. Section˜C.1 discusses error estimation, section˜C.2 uses the error estimate to implement SRC adaptively, and section˜C.3 optimizes the adaptive implementation using QR factorization updating algorithms.
@@ C.1.1 Leave-one-out error estimation: Matrix case
Surprisingly, the QR factorization ˜C.1 already contains enough information to compute an estimate of the error $\mathrm{Err}_{p}$. First, compute the inverse-adjoint of the matrix $R$:
@@
That is, $\smash{\widehat{\mathrm{Err}}}^{2}$ is a statistically unbiased estimate for the mean-square error $\mathrm{Err}_{p-1}^{2}$ of the rank-(p1)(p-1) randomized QB approximation. Consequently, $\smash{\widehat{\mathrm{Err}}}^{2}$ is an overestimate for $\mathrm{Err}_{p}^{2}$, on average.
@@ C.1.3 Norm estimation
This is an instance of the Girard–Hutchinson norm estimator, closely related to the Girard–Hutchinson trace estimator [40, §4]. In particular, we have the following standard guarantee [40, §4.8]:
@@
Under the hypotheses of Theorem˜4, the squared norm estimate $\smash{\widehat{\mathrm{Norm}}}^{2}$ is a statistically unbiased estimate for $\left\|\smash{B}^{(j)}\right\|^{2}_{\rm F}$. That is,
@@ C.2 Adaptive determination of bond dimension
Using the leave-one-out error estimate ˜C.2 and norm estimate ˜C.3, we can adaptively choose the dimension $\overline{\chi}$ adaptively for each bond in the output MPS using a tolerance. Let $\tau_{\rm abs}$ and $\tau_{\rm rel}$ be absolute and relative tolerances, and let $\overline{\chi}_{0}$ and $\Delta_{\chi}$ be a user-specified minimum bond dimension and increment. Begin each step $j$ of SRC using the minimum bond dimension $\overline{\chi}_{0}$, obtaining a $j$th site tensor $\eta^{(j)}$. Then, do the following steps:
@@
    Update randomized QB approximation. Compute the additional columns $\Delta_{\chi}$ of $Y^{(j)}$ using the (possibly updated) intermediate tensor $C^{(j-1)}$. Recompute (or update, see section˜C.3) the QR factorization of $Y^{(j)}$. Go to step 1.
When implemented using QR updating rather recomputation from scratch, as described in the next section, this adaptive version of the MPO–MPS multiplication algorithm has the same time complexity as reported in section˜3.5. In our experiments in section˜4.4.3, we use parameters $\overline{\chi}_{0}=2$, $\Delta_{\chi}=3$, and use only a relative tolerance (setting $\tau_{\rm abs}=0$). To implement with oversampling, we set the relative tolerance to be $0.1$ times the requested tolerance and run a final truncation with the requested tolerance.
@@ C.3 Final optimization: Updating the QR factorization
In the adaptive MPO–MPS multiplication algorithm (section˜C.2), we continue to accumulate “columns” in the 2- or 3-tensor $Y^{(j)}$ until a tolerance is satisfied. Recomputing a QR factorization each time we do this is computationally expensive, and we can make the algorithm more efficient by instead updating the QR factorization to include new columns. This material is standard but perhaps not well-known, so we provide a brief discussion.
@@
Another matrix that must be updated to implement the adaptive MPO–MPS multiplication algorithm efficiently is the matrix $G$ used in error estimation ˜C.2. Again using our superscript-less notation, we begin with $G=R^{-\dagger}$ and with to update $G$ to
@@
We recognize the original $G$ matrix in the top left corner of this matrix. To finish the formula for $G^{\prime}$, we just need to compute $(R^{\prime\prime})^{\dagger}$ and form the matrix triple product $((R^{\prime\prime})^{-\dagger}(R^{\prime})^{\dagger})G$. Since $R^{\prime\prime}$ is a small $\Delta_{\chi}\times\Delta_{\chi}$ matrix, this is efficient to do.
@@ Appendix D Full operation counts
Full operation counts for the methods in terms of all five parameters $n$, $d$, $\chi$, $D$, and $\overline{\chi}$ are presented in Table˜2. For simplified operation counts, see Table˜1.